If Amy gives 10 of her stamps to Robin, she will have four times as many stamps as Robin. If she gives 20 of her stamps to Robin, she will have three times as many stamps as Robin. How many stamps do they have in all?
Answer by Paul(988)

Let Amy's stamps be x
Let Robin's stamps be y
She loses 10 -->x-10
Other gains 10 -->y+10
4(y+10)=x-10
4y+40=x-10
x=4y+50 (subsitution)

She loses 20 -->x-20
other gains 20 -->y+20
3(y+20)=x-20
3y+60=x-20
3y+60=4y+50-20
y=30

x=4(30)+50
x=170
Hence, Amy has 170 stamps, and Robin has 30 stamps.
Paul.